### v3

# Limiting Sets for Digital Cones and Suspensions

Abstract
Cone and suspension constructions have been introduced in digital topology, modeled on those of classical topology. For digital cones and suspensions, and for some related digital images, we find $(m,n)$-limiting sets; especially $(0,0)$-limiting (freezing) sets.
Key words and phrases: digital topology, digital image, cone, suspension, freezing set, cold set, limiting set
MSC: 54B20, 54C35
1 Introduction
Although in some ways digitally continuous functions on digital images are similar to classical topology’s continuous functions on subsets of Euclidean spaces, the locally finite nature of digital images often restricts digitally continuous functions in ways that classical continuous functions are not restricted. The fixed point theory of digital images has yielded many such restrictions, leading to the study of freezing sets, $s$-cold sets, and their generalization, $(m,n)$-limiting sets.
We study freezing sets, $s$-cold sets, and $(m,n)$-limiting sets for digital cones, suspensions, and some cone-like and suspension-like digital images.
2 Preliminaries
We use ${\mathbb{N}}$ for the set of natural numbers, =${\mathbb{N}}^{*}={\mathbb{N}}\cup\{0\}$, ${\mathbb{Z}}$ for the set of integers, and #$\#X$ for the number of distinct members of $X$.
We typically denote a (binary) digital image as $(X,\kappa)$, where $X\subset{\mathbb{Z}}^{n}$ for some $n\in{\mathbb{N}}$ and $\kappa$ represents an adjacency relation of pairs of points in $X$. Thus, $(X,\kappa)$ is an undirected graph, in which members of $X$ may be thought of as black points, and members of ${\mathbb{Z}}^{n}\setminus X$ as white points, of a picture of some “real world” object or scene.
2.1 Adjacencies
Let $u,n\in{\mathbb{N}}$, $1\leq u\leq n$. The most used adjacencies in digital topology are the $c_{u}$ adjacencies, especially $c_{1}$ and $c_{n}$. These are defined as follows. Let $x,y\in{\mathbb{Z}}^{n}$,
$x=(x_{1},\ldots,x_{n}),~{}~{}~{}~{}~{}y=(y_{1},\ldots,y_{n}).$
We say $x$ and $y$ are $c_{u}$-adjacent if
  • $x\neq y$, and
  • for at most $u$ indices $i$, $\mid x_{i}-y_{i}\mid=1$, and
  • for all indices $j$ such that $\mid x_{j}-y_{j}\mid\neq 1$, we have $x_{j}=y_{j}$.
In low dimensions, many authors denote a $c_{u}$ adjacency by the number of points that can have this adjacency with a given point in ${\mathbb{Z}}^{n}$. E.g.,
  • In ${\mathbb{Z}}^{1}$, $c_{1}$-adjacency is 2-adjacency.
  • In ${\mathbb{Z}}^{2}$, $c_{1}$-adjacency is 4-adjacency and $c_{2}$-adjacency is 8-adjacency.
  • In ${\mathbb{Z}}^{3}$, $c_{1}$-adjacency is 8-adjacency, $c_{2}$-adjacency is 18-adjacency, and $c_{3}$-adjacency is 26-adjacency.
We use the notations $y\leftrightarrow_{\kappa}x$, or, when the adjacency $\kappa$ can be assumed, $y\leftrightarrow x$, to mean $x$ and $y$ are $\kappa$-adjacent. The notations $y\leftrightarroweq_{\kappa}x$, or, when $\kappa$ can be assumed, $y\leftrightarroweq x$, mean either $y=x$ or $y\leftrightarrow_{\kappa}x$.
Let
$N(X,x,\kappa)=\{\,y\in X\mid y\leftrightarroweq_{\kappa}x\,\}.$
A sequence $P=\{y_{i}\}_{i=0}^{m}$ in a digital image $(X,\kappa)$ is a $\kappa$-path from $a\in X$ to $b\in X$ if $a=y_{0}$, $b=y_{m}$, and $y_{i}\leftrightarroweq_{\kappa}y_{i+1}$ for $0\leq i<m$. We say $m$ is the length of the path:
$m=length(S).$
$X$ is $\kappa$-connected [12], or connected when $\kappa$ is understood, if for every pair of points $a,b\in X$ there exists a $\kappa$-path in $X$ from $a$ to $b$.
A (digital) $\kappa$-closed curve is a path $S=\{s_{i}\}_{i=0}^{m}$ such that $s_{0}=s_{m}$, and $0<|i-j|<m$ implies $s_{i}\neq s_{j}$. If, also, $0\leq i<m$ implies
$N(S,x_{i},\kappa)\setminus\{x_{i}\}=\{x_{(i-1)\,mod\,m},~{}x_{(i+1)\,mod\,m}\}$
then $S$ is a (digital) $\kappa$-simple closed curve.
Definition 2.1.
[9] Let $(X,\kappa)$ be a connected graph. The shortest path metric for $(X,\kappa)$ is
$d(x,y)=\min\{length(P)~{}|~{}P\mbox{ is a $\kappa$-path in $X$ from $x$ to $y$}\},\mbox{ for }x,y\in X.$
When it is useful to clarify the digital image considered, we use the notation $d_{(X,\kappa)}(x,y)$ for this metric.
2.2 Digitally continuous functions
Digital continuity preserves connectedness, as at Definition 2.2. By using adjacency as our notion of “closeness,” we get Theorem 2.3.
Definition 2.2.
[2] (generalizing a definition of [12]) Let $(X,\kappa)$ and $(Y,\lambda)$ be digital images. A function $f:X\rightarrow Y$ is $(\kappa,\lambda)$-continuous if for every $\kappa$-connected $A\subset X$ we have that $f(A)$ is a $\lambda$-connected subset of $Y$.
If $Y\subset X$, we use the abbreviation $\kappa$-continuous for $(\kappa,\kappa)$-continuous.
When the adjacency relations are understood, we will simply say that $f$ is continuous. Continuity can be expressed in terms of adjacency of points:
Theorem 2.3.
[12, 2] A function $f:X\to Y$ is $(\kappa,\lambda)$-continuous if and only if $x\leftrightarrow_{\kappa}x^{\prime}$ implies $f(x)\leftrightarroweq_{\lambda}f(x^{\prime})$.
See also [7, 8], where similar notions are referred to as immersions, gradually varied operators, and gradually varied mappings.
A digital isomorphism (called homeomorphism in [1]) is a $(\kappa,\lambda)$-continuous surjection $f:X\to Y$ such that $f^{-1}:Y\to X$ is $(\lambda,\kappa)$-continuous.
The literature uses path polymorphically: a $(c_{1},\kappa)$-continuous function $f:[0,m]_{{\mathbb{Z}}}\to X$ is a $\kappa$-path if $f([0,m]_{{\mathbb{Z}}})$ is a $\kappa$-path from $f(0)$ to $f(m)$ as described above.
We define [6]
[…]
Let $X\subset{\mathbb{Z}}^{n}$. The projection functions $p_{i}:X\to{\mathbb{Z}}$ are defined by
$p_{i}(x_{1},\ldots,x_{i},\ldots,x_{n})=x_{i}.$
The projection functions are $(c_{u},c_{1})$-continuous for $1\leq u\leq n$ [9].
For a function $f:X\to X$, we denote by $\operatorname{Fix}(f)$ the set of fixed points of $f$, i.e., $x\in\operatorname{Fix}(f)$ means $f(x)=x$.
2.3 Freezing, cold, and limiting sets
Material in this section is largely quoted or paraphrased from [4, 5].
In a Euclidean space, knowledge of the fixed point set of a continuous self-map $f:X\to X$ often gives little information about $f|_{X\setminus\operatorname{Fix}(f)}$. By contrast, knowledge of $\operatorname{Fix}(f)$ for $f\in C(X,\kappa)$ can tell us much about $f|_{X\setminus\operatorname{Fix}(f)}$. This motivates the study of freezing and cold sets.
Definition 2.4.
[4] Let $(X,\kappa)$ be a digital image. We say $A\subset X$ is a freezing set for $(X,\kappa)$ if given $g\in C(X,\kappa)$, $A\subset\operatorname{Fix}(g)$ implies $g=\operatorname{id}_{X}$. If no proper subset of a freezing set $A$ is a freezing set for $(X,\kappa)$, then $A$ is a minimal freezing set for $(X,\kappa)$.
Theorem 2.5.
[4] Let $X=\Pi_{i=1}^{n}[0,m_{i}]_{{\mathbb{Z}}}$. Let $A=\Pi_{i=1}^{n}\{0,m_{i}\}$.
  • Let $Y=\Pi_{i=1}^{n}[a_{i},b_{i}]_{{\mathbb{Z}}}$ be such that $X\subset Y$. Let $f:X\to Y$ be $c_{1}$-continuous. If $A\subset\operatorname{Fix}(f)$, then $X\subset\operatorname{Fix}(f)$.
  • $A$ is a freezing set for $(X,c_{1})$; minimal for $n\in\{1,2\}$.
Theorem 2.6.
[4] Let $A$ be a freezing set for the digital image $(X,\kappa)$ and let $F:(X,\kappa)\to(Y,\lambda)$ be an isomorphism. Then $F(A)$ is a freezing set for $(Y,\lambda)$.
The following are useful for determining fixed point and freezing sets.
Proposition 2.7.
(Corollary 8.4 of [6]) Let $(X,\kappa)$ be a digital image and $f\in C(X,\kappa)$. Suppose $x,x^{\prime}\in\operatorname{Fix}(f)$ are such that there is a unique shortest $\kappa$-path $P$ in $X$ from $x$ to $x^{\prime}$. Then $P\subseteq\operatorname{Fix}(f)$.
Lemma 2.8, below,
$\ldots$ can be interpreted to say that in a $c_{u}$-adjacency, a continuous function that moves a point $p$ also moves a point that is “behind” $p$. E.g., in ${\mathbb{Z}}^{2}$, if $q$ and $q^{\prime}$ are $c_{1}$- or $c_{2}$-adjacent with $q$ left, right, above, or below $q^{\prime}$, and a continuous function $f$ moves $q$ to the left, right, higher, or lower, respectively, then $f$ also moves $q^{\prime}$ to the left, right, higher, or lower, respectively [4].
Lemma 2.8.
[4] Let $(X,c_{u})\subset{\mathbb{Z}}^{n}$ be a digital image, $1\leq u\leq n$. Let $q,q^{\prime}\in X$ be such that $q\leftrightarrow_{c_{u}}q^{\prime}$. Let $f\in C(X,c_{u})$.
  1. 1.
    If $p_{i}(f(q))>p_{i}(q)>p_{i}(q^{\prime})$ then $p_{i}(f(q^{\prime}))>p_{i}(q^{\prime})$.
  2. 2.
    If $p_{i}(f(q))<p_{i}(q)<p_{i}(q^{\prime})$ then $p_{i}(f(q^{\prime}))<p_{i}(q^{\prime})$.
Definition 2.9.
[11] Let $X\subset{\mathbb{Z}}^{n}$. The boundary of $X$ is
$Bd(X)=\{x\in X\,|\mbox{ there exists }y\in{\mathbb{Z}}^{n}\setminus X\mbox{ such that }y\leftrightarrow_{c_{1}}x\}.$
Theorem 2.10.
[4] Let $X\subset{\mathbb{Z}}^{n}$ be finite. Then for $1\leq u\leq n$, $Bd(X)$ is a freezing set for $(X,c_{u})$.
In particular, we have the following.
Theorem 2.11.
[4] Let $X=\prod_{i=1}^{n}[0,m_{i}]_{{\mathbb{Z}}}\subset{\mathbb{Z}}^{n}$, where $m_{i}>1$ for all $i$. Then $Bd(X)$ is a minimal freezing set for $(X,c_{n})$.
Definition 2.12.
[4] Given $s\in{\mathbb{N}}^{*}$, we say $A\subset X$ is an $s$-cold set for the connected digital image $(X,\kappa)$ if given $f\in C(X,\kappa)$ such that $f|_{A}=\operatorname{id}_{A}$, then for all $x\in X$, $d_{(X,\kappa)}(x,f(x))\leq s$. If no proper subset of $A$ is an $s$-cold set for $(X,\kappa)$, then $A$ is minimal. A cold set is a 1-cold set.
Theorem 2.13.
[4] Let $(X,\kappa)$ be a connected digital image, let $A$ be an $s$-cold set for $(X,\kappa)$, and let $F:(X,\kappa)\to(Y,\lambda)$ be an isomorphism. Then $F(A)$ is an $s$-cold set for $(Y,\lambda)$.
Definition 2.14.
[5] Let $X$ be a $\kappa$-connected digital image. Let $f\in C(X,\kappa)$ and let $A\subset X$. Let $m,n\in{\mathbb{N}}^{*}$. If for all $x\in A$ we have $d_{(X,\kappa)}(x,f(x))\leq n$, we say $f|_{A}$ is an $n$-map. We say $f$ is an $n$-map if $f|_{X}$ is an $n$-map. If for all $f\in C(X,\kappa)$, $f|_{A}$ being an $m$-map implies $f$ is an $n$-map, then $A$ is an $(m,n)$-limiting set for $(X,\kappa)$ and $(X,\kappa)$ is $(A,m,n)$-limited. Such a set is a minimal $(m,n)$-limiting set for $(X,\kappa)$ if no proper subset $A^{\prime}$ of $A$ is an $(m,n)$-limiting set for $(X,\kappa)$.
2.4 Cones and suspensions
The paper [10] introduces cone and suspension constructions in digital topology, according to the following.
Definition 2.15.
Let $(X,\kappa)$ be a digital image.
  • The cone on $X$ is the set $CX=X\cup\{U\}$, where $U\not\in X$, and adjacency $\kappa^{\prime}$, where $U\leftrightarrow_{\kappa^{\prime}}x$ for all $x\in X$ and if $a,b\in X$ then $a\leftrightarrow_{\kappa^{\prime}}b$ if and only if $a\leftrightarrow_{\kappa}b$.
  • The suspension on $X$ is the set $SX=X\cup\{U,L\}$, where $U\neq L$ and $\{U,L\}\cap X=\emptyset$; and adjacency $\kappa^{\prime}$, where $U\leftrightarrow_{\kappa^{\prime}}x\leftrightarrow_{\kappa^{\prime}}L$ for all $x\in X$; $U$ and $L$ are not adjacent; and if $a,b\in X$ then $a\leftrightarrow_{\kappa^{\prime}}b$ if and only if $a\leftrightarrow_{\kappa}b$.
One may consider $U$ as the “upper pole” of $CX$ and of $SX$, and $L$ as the “lower pole” of $SX$.
When $\kappa$ can be assumed, we often abbreviate $(CX,\kappa^{\prime})$ as $CX$, and $(SX,\kappa^{\prime})$ as $SX$.
For both $(CX,\kappa^{\prime})$ and $(SX,\kappa^{\prime})$, we implicitly use the following.
Theorem 2.16.
[10] Any finite graph $(X,\kappa)$ can be isomorphically mapped onto a subset of $(Z^{n-1},c_{n-1})$, where $n=\#X$.
3 Limiting sets for $CX$ and $SX$
Remark 3.1.
$d_{SX}(U,L)=2$.
Given a graph $(X,\kappa)$ and $D\subset X$, we say $D$ is a $\kappa$-dominating set in $X$ if for every $x\in X$ there exists $d\in D$ such that $d\leftrightarroweq_{\kappa}x$.
Proposition 3.2.
Let $(X,\kappa)$ be a digital image. Let $f\in C(X,\kappa)$. Let $D\subset X$ such that $D$ is a $\kappa$-dominating set in $X$. If $f|_{D}$ is an $m$-map, then $f$ is an $(m+2)$-map.
Proof.
Assume $D$ is a $\kappa$-dominating set in $X$. For each $x\in X$, there exists $\delta_{x}\in D$ such that $x\leftrightarroweq_{\kappa}\delta_{x}$. By continuity
$d_{(X,\kappa)}(x,f(x))\leq d(x,\delta_{X})+d(\delta_{X},f(\delta_{X}))+d(f(\delta_{X}),f(x))$ (1)
$\leq 1+m+1=m+2.$
The assertion follows. ∎
Proposition 3.3.
Let $(X,\kappa)$ be a digital image. Given $x,y\in CX$, $a,b\in SX$, we have $d_{CX}(x,y)\leq 2$, $d_{SX}(a,b)\leq 2$.
Proof.
We must show that there are paths of length at most 2 from $x$ to $y$ in $CX$ and from $a$ to $b$ in $SX$.
Assume without loss of generality that $x\neq y$. For $\{x,y\}\subset CX$, $\{x,U,y\}$ is a path of length at most 2.
In $SX$, for $\{a,b\}\subset X$, $\{a,U,b\}$ is a path in $SX$ from $a$ to $b$ of length 2. For $a\in X$, $\{a,L\}$ is a path in $SX$ of length 1 from $a$ to $L$. For $b\in X$, $\{U,b\}$ is a path in $SX$ of length 1 from $U$ to $b$. For any $x\in X$, $\{U,x,L\}$ is a path in $SX$ from $U$ to $L$ of length 2. The assertion follows. ∎
Remark 3.4.
[…]
Theorem 3.5.
Let $(X,\kappa)$ be a digital image such that
for every $x\in X$, $N(X,x,\kappa)\neq X$ (2)
Then $X$ is a minimal freezing set for $CX$.
Proof.
Suppose $x_{0}\in X$. Then the function $f:CX\to CX$ given by
$f(x)=\left\{\begin{array}[]{ll}U&\mbox{if }x=x_{0},\\ x&\mbox{if }x\neq x_{0},\end{array}\right.$
is easily seen to be continuous, with $f|_{CX\setminus\{x_{0}\}}=\operatorname{id}_{CX\setminus\{x_{0}\}}$ but $f\neq\operatorname{id}_{CX}$. Thus we must have $x_{0}$ belonging to every freezing set for $CX$. Therefore
$X\mbox{ is a subset of every freezing set for }CX.$ (3)
The function $f:CX\to CX$ given, for any $y\in X$, by
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\in X,\\ y&\mbox{if }x=U,\end{array}\right.$
cannot be continuous, by (2). Therefore, if a function $g:CX\to CX$ is continuous and $g|_{X}=\operatorname{id}_{X}$, we must have $g(U)=U$, so $g=\operatorname{id}_{CX}$. Therefore, $X$ is a freezing set for $CX$. It follows from (3) that $X$ is minimal. ∎
Proposition 3.6.
Let $B$ be a freezing set for $SX$. Then $\{\,U,L\,\}\subset B$.
Proof.
If $U\not\in B$, then the function $f:SX\to SX$ given by
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U,\\ L&\mbox{if }x=U\end{array}\right.$
is easily seen to be continuous, satisfies $f|_{B}=\operatorname{id}_{B}$, and $f\neq\operatorname{id}_{SX}$. Thus we must have $U\in B$.
Similarly, if $L\not\in B$, then the function $g:SX\to SX$ given by
$g(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq L,\\ U&\mbox{if }x=L,\end{array}\right.$
is easily seen to be a continuous function that satisfies $g|_{B}=\operatorname{id}_{B}$, and $g\neq\operatorname{id}_{SX}$. Thus we must have $L\in B$. ∎
Lemma 3.7.
Let $f\in C(SX,\kappa^{\prime})$ such that $f|_{\{\,U,L\,\}}=\operatorname{id}_{\{\,U,L\,\}}$. Then $f(X)\subset X$.
Proof.
Suppose otherwise. Then for some $x\in X$, either $f(x)=U$ or $f(x)=L$. In the former case, we have $x\leftrightarrow L$; but $f(x)=U$ and $f(L)=L$ are neither adjacent nor equal, contrary to the continuity of $f$. In the latter case, $x\leftrightarrow U$; but $f(x)=L$ and $f(U)=U$ are neither adjacent nor equal, contrary to the continuity of $f$. These contradictions establish the assertion. ∎
Theorem 3.8.
Let $A\subset(X,\kappa)$ and $B=A\cup\{\,U,L\,\}\subset SX$. Then $A$ is a (minimal) freezing set for $(X,\kappa)$ if and only if $B$ is a (minimal) freezing set for $SX$.
Proof.
Suppose $A$ is a freezing set for $(X,\kappa)$. Let $f:SX\to SX$ be continuous, such that $f|_{B}=\operatorname{id}|_{B}$. By Lemma 3.7, $f|_{X}:X\to X$; and $f|_{A}=\operatorname{id}_{A}$. By choice of $A$, $f|_{X}=\operatorname{id}_{X}$. Thus $f=\operatorname{id}_{SX}$, so $B$ is freezing.
Suppose $A$ is a minimal freezing set for $(X,\kappa)$. Let $B^{\prime}\subset B$ such that $B^{\prime}$ is a freezing set for $(SX,\kappa^{\prime})$. Let $A^{\prime}=B^{\prime}\setminus\{U,L\}\subset A$. Suppose $f\in C(X,\kappa)$ such that $f|_{A^{\prime}}=\operatorname{id}_{A^{\prime}}$. Let $g:SX\to SX$ be defined by
$g(y)=\left\{\begin{array}[]{ll}y&\mbox{if }y\in A^{\prime}\cup\{U,L\},\\ f(y)&\mbox{otherwise.}\end{array}\right.$
We see easily that $g\in C(SX,\kappa^{\prime})$ and $g|_{B^{\prime}}=\operatorname{id}_{B^{\prime}}$. By choice of $B^{\prime}$, $g=\operatorname{id}_{SX}$. Therefore, $f=\operatorname{id}_{X}$. Thus $A^{\prime}$ is a freezing set for $(X,\kappa)$. Since $A^{\prime}\subset A$ and $A$ is minimal, we have $A^{\prime}=A$, so by Proposition 3.6, $B^{\prime}=B$. Thus $B$ is a minimal freezing set for $(SX,\kappa^{\prime})$.
Let $B$ be a freezing set for $SX$. By Proposition (3.6),
$\{\,U,L\,\}\subset B$ (4)
Let $f\in C(X,\kappa)$ such that $f|_{A}=\operatorname{id}_{A}$. Then the function $g:SX\to SX$ given by
$g(y)=\left\{\begin{array}[]{ll}f(y)&\mbox{if }y\in X,\\ y&\mbox{if }y\in\{U,L\},\end{array}\right.$
is easily seen to belong to $C(SX,\kappa^{\prime})$, with $g|_{B}=\operatorname{id}_{B}$. By choice of $B$, $g=\operatorname{id}_{CX}$. Therefore, $f=g|_{X}=\operatorname{id}_{X}$. Thus $A$ is a freezing set for $(X,\kappa)$.
If a subset $A^{\prime}$ of $A$ is also a freezing set for $X$, by the above and Proposition 3.6, $B^{\prime}=A^{\prime}\cup\{U,L\}$ is a freezing set for $CX$. Since $B^{\prime}\subset B$ and $B$ is minimal for $CX$, we must have $B^{\prime}=B$, hence $A^{\prime}=A$. Hence $A$ is minimal for $X$. ∎
Theorem 3.9.
Let $(X,\kappa)$ be a digital image that has the property (2). Then $U$ is not a member of a minimal cold set of $CX$; and neither $U$ nor $L$ is a member of a minimal cold set of $SX$.
Proof.
Let $Y$ be a cold set for $(CX,\kappa^{\prime})$. Since every member of  $CX$ is $\kappa^{\prime}$-adjacent or equal to $U$, it follows that $Y\setminus\{U\}$ is a cold set for $(CX,\kappa^{\prime})$. Thus $U$ is not a member of a minimal cold set for $(CX,\kappa^{\prime})$.
Let $Y$ be a cold set for $(SX,\kappa^{\prime})$. Let $f\in C(SX,\kappa^{\prime})$ be such that $f|_{Y}=\operatorname{id}_{Y}$. Then $f$ is a 1-map, so $f(U)\neq L$ and $f(L)\neq U$. By reasoning similar to the above, neither $U$ nor $L$ is a member of a minimal cold set of $SX$. ∎
Theorem 3.10.
Let $(X,\kappa)$ be a digital image. Then $U$ is not a member of a minimal (1,1)-limiting subset $A$ of $CX$.
Proof.
Suppose $U\in A$ where $A$ is (1,1)-limiting for $CX$. Let $A^{\prime}=A\setminus\{U\}$. Let $f\in C(CX,\kappa^{\prime})$ such that $f|_{A^{\prime}}$ is a 1-map. Since $f|_{\{U\}}$ must be a 1-map, it follows that $f|_{A}$ is a 1-map, so by choice of $A$, $f$ is a 1-map. Thus $A^{\prime}$ is (1,1)-limiting for $CX$; therefore $A$ is not minimal. ∎
Theorem 3.11.
Let $(X,\kappa)$ be a digital image with property (2). Then $U$ and $L$ are members of every $A\subset SX$ such that $SX$ is $(A,1,1)$-limited.
Proof.
Let $A\subset X$ be such that
$SX\mbox{ is }(A,1,1)-\mbox{limited.}$. (5)
We argue by contradiction. Suppose $U\not\in A$. Let $f:SX\to SX$ be the function
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq U,\\ L&\mbox{if }x=U.\end{array}\right.$
It is easily seen that $f\in C(SX,\kappa^{\prime})$ and $f|_{A}$ is a 1-map, but by Remark 3.1, $d_{SX}(U,f(U))=2$, contrary to (5). Thus we must have $U\in A$.
A similar argument shows we must have $L\in A$. ∎
4 Freezing sets for digital pyramids
In this section, we consider freezing sets for other cone-like digital images. In particular, for $n\in{\mathbb{N}}$, we consider a digital version of the surface of a pyramid (see Figures 1 and 2)
$P_{n}=\bigcup_{i=0}^{n}T_{i}\subset{\mathbb{Z}}^{3}$
where
$T_{i}=(\{\,-i,i\,\}\times[-i,i]_{{\mathbb{Z}}}\times\{n-i\})\cup([-i,i]_{{\mathbb{Z}}}\times\{\,-i,i\,\}\times\{n-i\}).$
Also, let
$T_{i}^{\prime}=\{\,(-i,-i,n-i),(i,-i,n-i),(i,i,n-i),(-i,i,n-i)\,\}\subset T_{i},$
$W_{i}=[-i,i]_{{\mathbb{Z}}}^{2}\times\{n-i\}.$
Figure 2 illustrates that in the $c_{3}$ adjacency, each $T_{i}$ is a square simple closed curve whose “corners” are the members of $T_{i}^{\prime}$; the two $c_{1}$-neighbors in $T_{i}$ of each “corner” are $c_{2}$-adjacent. $P_{n}$ can be thought of as a digital analog of a cone on a simple closed curve - the latter represented by $T_{n}$.
Let
$Q_{n}=\bigcup_{i=0}^{n}W_{i}\subset{\mathbb{Z}}^{3},$ (6)
which can be thought of as a digital analog of a solid tetrahedron.
For each of $P_{n},Q_{n}$, we let $U=(0,0,n)$, the high point of the image.
Figure 1: Left and front sides of the digital pyramid $(P_{2},c_{3})$ with base $[-2,2]_{{\mathbb{Z}}}^{2}$.
Marked edges:
LR (left rear) from (-2,-2,0) to U
LF (left front) from (2,-2,0) to U
RF (right front) from (2,2,0) to U
Faces shown:
L (left) - the digital triangle with vertices $U$, $(-2,-2,0)$, and $(2,-2,0)$
F (front) - the digital triangle with vertices $U$, $(2,-2,0)$, and $(2,2,0)$
Figure 2: Slices of $(P_{n},c_{3})$
(a) an isomorphic copy in $({\mathbb{Z}}^{2},c_{2})$ of $(T_{1},c_{3})$
(b) an isomorphic copy in $({\mathbb{Z}}^{2},c_{2})$ of $(T_{2},c_{3})$
We will refer to the edges of the pyramid containing $U$ as follows.
$LR$ (left rear): the digital segment from $(-n,-n,0)$ to $U$
$LF$ (left front): the digital segment from $(n,-n,0)$ to $U$
$RF$ (right front): the digital segment from $(n,n,0)$ to $U$
$RR$ (right rear): the digital segment from $(-n,n,0)$ to $U$
$BL$ (bottom left): the digital segment from $(-n,-n,0)$ to $(n,-n,0)$
$BF$ (bottom front): the digital segment from $(n,-n,0)$ to $(n,n,0)$
$BR$ (bottom right): the digital segment from $(n,n,0)$ to $(-n,n,0)$
$BB$ (bottom back): the digital segment from $(-n,n,0)$ to $(-n,-n,0)$
Faces of the pyramid are the digital triangles as follows:
$L$ (left): edges are $BL,LR,LF$
$F$ (front): edges are $BF,LF,RF$
$R$ (right): edges are $BR,RF,RB$
$B$ (back): edges are $BB,RB,LB$
4.1 Freezing sets for $P_{n}$
It is easily seen (although perhaps easily overlooked) that $(P_{n},c_{1})$ is not connected, as each $T_{i}$ is a $c_{1}$-component. Therefore, we consider $(P_{n},c_{3})$.
Proposition 4.1.
Let $A$ be a freezing set for $(P_{n},c_{3})$. Then $T_{n}\subset A$.
Proof.
We will show that for all $x\in T_{n}$ there exists $f\in C(P_{n},c_{3})$ such that
$f|_{P_{n}\setminus\{x\}}=\operatorname{id}_{P_{n}\setminus\{x\}}\mbox{ and }f(x)\neq x.$ (7)
The assertion will follow.
For the corner point $x=(n,-n,0)$, the function
$f(p)=\left\{\begin{array}[]{ll}(n-1,-n+1,1)&\mbox{if }p=x,\\ p&\mbox{if }p\neq x\end{array}\right.$
is easily seen to belong to $C(P_{n},c_{3})$ and to satisfy (7). See Figure 1, using $n=2$, for illustration. We can similarly construct, for the other corner points $(-n,-n,0)$, $(n,n,0)$, and $(-n,n,0)$ of $T_{n}$, a function $f\in C(P_{n},c_{3})$ that satisfies (7).
A point of $BF$ that is not a corner of $P_{n}$ must be of the form $x=(n,z,0)$, where $-n<z<n$. The function
$f(p)=\left\{\begin{array}[]{ll}(n-1,z,1)&\mbox{if }p=x,\\ p&\mbox{if }p\neq x\end{array}\right.$
is easily seen to belong to $C(P_{n},c_{3})$ and to satisfy (7). See Figure 1, using $n=2$ and any $z\in\{\,-1,0,1\,\}$, for illustration. We can similarly construct, for the other non-corner points of $T_{n}$, a function $f\in C(P_{n},c_{3})$ that satisfies (7). ∎
Theorem 4.2.
$T_{n}$ is the only minimal freezing set for $P_{n}$.
Proof.
The function $F:P_{n}\to[-n,n]_{{\mathbb{Z}}}^{2}$ given by $F(a,b,c)=(a,b)$ is a $(c_{3},c_{2})$-isomorphism. This can be seen by imagining looking down on $P_{n}$ from above $U$. Since
$F(T_{n})=([-n,n]_{{\mathbb{Z}}}\times\{-n,n\})\cup(\{-n,n\}\times[-n,n]_{{\mathbb{Z}}})=Bd([-n,n]_{{\mathbb{Z}}}^{2})$
when $[-n,n]_{{\mathbb{Z}}}^{2}$ is considered as a subset of ${\mathbb{Z}}^{2}$, it follows from Theorem 2.6 and Theorem 2.11 that $T_{n}$ is a minimal freezing set for $(P_{n},c_{3})$.
The uniqueness of $T_{n}$ as a minimal freezing set for $(P_{n},c_{3})$ follows from Proposition 4.1. ∎
4.2 Freezing sets for digital solid pyramids
In this section, we consider digital images $(Q_{n},c_{3})$, where $Q_{n}$ is as defined at (6). Notice $Q_{n}$ is a cone-like structure over a base of the (solid) digital square $W_{n}$; $Q_{n}$ may be regarded as a solid version of $P_{n}$.
As above, we let $U$ be the highest point of $Q_{n}$: $U=(0,0,n)$.
We have the following.
Proposition 4.3.
$\{U\}\cup W_{n}$ is a subset of every freezing set for $(Q_{n},c_{3})$.
Proof.
Let $A\subset Q_{n}$. Suppose $U\not\in A$. Then the function $f:Q_{n}\to Q_{n}$ given by
$f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (0,0,n-1)&\mbox{if }x=U,\end{array}\right.$
is easily seen to belong to $C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{Q_{n}}$. Thus it is necessary for $U\in A$ in order for $A$ to be freezing.
Let $y\in W_{n}$. We show that if $y\not\in A$ then $A$ is not a freezing set for $(Q_{n},c_{3})$.
  • If $y=(a,b,0)$ is in the “interior” of $W_{n}$, i.e., if $y\in[n-1,n-1]_{{\mathbb{Z}}}^{2}\times\{0\}$, the function $f:Q_{n}\to Q_{n}$ given by
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (a,b,1)&\mbox{if }x=y,\end{array}\right.$
    is easily seen to belong to $C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{Q_{n}}$.
  • If $y=(-n,b,0)\in BB$ then the function $f:Q_{n}\to Q_{n}$ given by
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (-(n-1),b,0)&\mbox{if }x=y,\end{array}\right.$
    is easily seen to belong to $C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{Q_{n}}$.
  • If $y=(n,b,0)\in BF$ then the function $f:Q_{n}\to Q_{n}$ given by
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (n-1,b,0)&\mbox{if }x=y,\end{array}\right.$
    is easily seen to belong to $C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{Q_{n}}$.
  • If $y=(a,-n,0)\in BL$ then the function $f:Q_{n}\to Q_{n}$ given by
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (a,-(n-1),0)&\mbox{if }x=y,\end{array}\right.$
    is easily seen to belong to $C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{Q_{n}}$.
  • If $y=(a,n,0)\in BR$ then the function $f:Q_{n}\to Q_{n}$ given by
    $f(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\neq y,\\ (a,n-1,0)&\mbox{if }x=y,\end{array}\right.$
    is easily seen to belong to $C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$ and $f\neq\operatorname{id}_{Q_{n}}$.
Thus, in all cases, we need $x$ as a member of $A$ in order for $A$ to be freezing. ∎
Theorem 4.4.
$A=\{U\}\cup W_{n}$ is a minimal freezing set for $(Q_{n},c_{3})$.
Proof.
By Proposition 4.3, it suffices to show that $A$ is freezing. Let $f\in C(Q_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$.
We show
$p_{3}(f(x))=p_{3}(x)\mbox{ for all }x\in Q_{n}.$ (8)
For all $x\in Q_{n}$ there is a $c_{3}$-path $P=\{x_{i}\}_{i=0}^{n}$ from a point $x_{0}\in W_{n}$ to $U$ such that $x\in P$ and $p_{3}(x_{i})=i$ for $0\leq i\leq n$. It follows from Lemma 2.8 that if $p_{3}(f(x))<p_{3}(x)$, then pulling points of $P$ to lesser $3^{rd}$ coordinates would propagate so that $p_{3}(f(U))<p_{3}(U)$, contrary to $U\in\operatorname{Fix}(f)$; and if $p_{3}(f(x))>p_{3}(x)$, then pulling points of $P$ to greater $3^{rd}$ coordinates would propagate so that $p_{3}(f(x_{0}))>p_{3}(x_{0})$, contrary to $x_{0}\in W_{n}\subset\operatorname{Fix}(f)$. Thus, (8) is established.
Next, we show
$LR\cup LF\cup RF\cup RR\subset\operatorname{Fix}(f).$ (9)
To establish the latter claim, we will show $LF\subset\operatorname{Fix}(f)$. Notice $LF$ is the unique shortest $c_{3}$-path in $Q_{n}$ from $(n,-n,0)\in A$ to $U\in A$. By Proposition 2.7, $LF\subset\operatorname{Fix}(f)$ A similar argument shows that $LR\cup RF\cup RR\subset\operatorname{Fix}(f)$, and (9) follows.
Now, we claim
$\begin{array}[]{l}u\in L\cup R\mbox{ implies }p_{1}(f(u))=p_{1}(u);\\ v\in F\cup B\mbox{ implies }p_{2}(f(v))=p_{2}(v).\end{array}$ (10)
Suppose $u=(a,b,c)\in L$. Then $u\in\{(i,b,c)\}_{i=-(n-c)}^{(n-c)}$, a $c_{3}$-path from $u_{0}=(-(n-c),b,n-c)\in LR\subset\operatorname{Fix}(f)$ to $u^{\prime}=(n-c,b,n-c)\in LF\subset\operatorname{Fix}(f)$. If $p_{1}(f(u))>p_{1}(u)$ then by Lemma 2.8 we would have $p_{1}(f(u_{0}))>p_{1}(u_{0})$, contrary to $u_{0}\in\operatorname{Fix}(f)$. If $p_{1}(f(u))<p_{1}(u)$ then by Lemma 2.8 we would have $p_{1}(f(u^{\prime}))<p_{1}(u^{\prime})$, contrary to $u^{\prime}\in\operatorname{Fix}(f)$. Hence $p_{1}(f(u))=p_{1}(u)$. We show the other claims of (10) similarly.
Now let $q=(a,b,c)\in Q_{n}$. Then $q$ belongs to the paths
$P_{1}=\{(i,b,c)\}_{i=-(n-c)}^{n-c},~{}~{}~{}~{}~{}P_{2}=\{(a,j,c)\}_{j=-(n-c)}^{n-c}$
with endpoints in $B\cup F$ and in $L\cup R$, respectively. Note the sequence of points in $P_{1}$ is monotone increasing in the first coordinate. It follows as above from (10) and Lemma 2.8 that we must have $p_{1}(f(q))=p_{1}(q)$. Similarly, $p_{2}(f(q))=p_{2}(q)$.
From (8), it follows that $f(q)=q$, i.e., $f=\operatorname{id}_{Q_{n}}$. Hence $A$ is freezing. ∎
5 Freezing sets for suspension-like images
Let
$H_{n}=P_{n}\cup P_{n}^{\prime},\mbox{ where }P_{n}^{\prime}=\{\,(a,b,c)\in{\mathbb{Z}}^{3}\mid(a,b,-c)\in P_{n}\,\}$
$H_{n}$ may be regarded as a digital model of the topological suspension of the square (border only) modeled by $T_{n}$. $H_{n}$ may also be regarded as made of 2 copies of $P_{n}$ with their bases sewn together.
Let
$K_{n}=Q_{n}\cup Q_{n}^{\prime},\mbox{ where }Q_{n}^{\prime}=\{\,(a,b,c)\in{\mathbb{Z}}^{3}\mid(a,b,-c)\in Q_{n}\,\}$
$K_{n}$ may be regarded as a digital model of the topological suspension of the square (with interior) modeled by $W_{n}$. $K_{n}$ may also be regarded as made of 2 copies of $Q_{n}$ with their bases sewn together.
As above, we let $U$ be the highest point, and $L$, the lowest point, of $H_{n}$ and of $K_{n}$:
$U=(0,0,n),~{}~{}~{}~{}~{}L=(0,0,-n).$
See Figure 3.
Figure 3: Left and front sides of the suspension-like digital image $(H_{2},c_{3})$
We have the following.
Theorem 5.1.
$T_{n}$ is a freezing set for $(H_{n},c_{3})$.
Proof.
We show $T_{n}$ is freezing, as follows. Let $f\in C(H_{n},c_{3})$ such that $f|_{T_{n}}=\operatorname{id}_{T_{n}}$.
First, we show $f(P_{n})\subset P_{n}$. Let $x=(a,b,c)\in P_{n}$. Suppose, in order to obtain a contradiction, that $f(x)\not\in P_{n}$. Then $p_{3}(f(x))<0$, so
$d(x,U)=n-c<n+\mid p_{3}(f(x))\mid\,\leq d(f(x),U))=d(f(x),f(U)),$
which contradicts the continuity of $f$. Thus $f(P_{n})\subset P_{n}$.
Thus $f|_{P_{n}}:P_{n}\to P_{n}$ satisfies $f|_{T_{n}}=\operatorname{id}_{T_{n}}$. From Theorem 4.2 it follows that $f|_{P_{n}}=\operatorname{id}_{P_{n}}$.
Similarly, $f|_{P_{n}^{\prime}}=\operatorname{id}_{P_{n}^{\prime}}$. It follows that $T_{n}$ is a freezing set for $(H_{n},c_{3})$. ∎
Proposition 5.2.
$\{U,L\}\cup T_{n}$ is a subset of every freezing set for $(K_{n},c_{3})$.
Proof.
The assertion follows as in the proof of Proposition 4.3. ∎
Theorem 5.3.
$\{U,L\}\cup T_{n}$ is a minimal freezing set for $(K_{n},c_{3})$.
Proof.
In light of Proposition 5.2, it suffices to show that $A=\{U,L\}\cup T_{n}$ is a freezing set for $(K_{n},c_{3})$. Let $f\in C(K_{n},c_{3})$ such that $f|_{A}=\operatorname{id}_{A}$.
Given $x\in K_{n}\setminus\{U,L\}$, we have as in the proof of Theorem 4.4 that
for $\mbox{for all }x\in K_{n},~{}~{}~{}p_{3}(f(x))=p_{3}(x).$ (11)
In particular, $f(Q_{n})\subset Q_{n}$. Similarly, $f(Q_{n}^{\prime})\subset Q_{n}^{\prime}$.
The assertion follows as in Theorem 4.4. ∎
6 Further remarks
Models for digital topology of the topological constructions of cone and suspension were developed by Lupton, Oprea, and Scoville in [10]. We have obtained some $(m,n)$-limiting sets for these constructions. We have also found freezing sets for some cone-like and suspension-like digital images.